OpenStudy (anonymous):

derivate of (3X)* ((sinx / X - X^2 / 3))

OpenStudy (bahrom7893):

y = (3X)* ((sinx / X - X^2 / 3)) = 3Sin(x) - x^3/3 y' = 3Cos(x) - 3x^2/3 = 3Cos(x) - x^2

OpenStudy (bahrom7893):

become a fan if I helped, thanks!